Vines Corporation produces custom machine parts on a job order basis. The company has two direct product cost categories: direct materials and direct labour. In the past, indirect manufacturing costs were allocated to products using a single indirect cost pool, allocated based on direct labour hours. The indirect cost rate was $100 per direct labour hour.

The managers of Vines Corporation decided to switch from a manual system to software programs that release materials and that signal machines when to begin working. Simultaneously, the company adopted an activity-based costing system. The manufacturing process has been organized into six activities, each with its own supervisor who is responsible for controlling costs. The following list indicates the activities, cost drivers, and cost allocation rates:

Cost Cost per Unit
Activity Driver of Cost Driver
Material handling Number of parts $ 0.40
Milling Machine hours 36
Grinding Number of parts 0.70
Assembly Hours spent in assembly 5
Inspection Number of units produced 30
Shipping Number of orders shipped 1,600

The company’s information system automatically collects the necessary data for these six activity areas. The data for two recent jobs follow:

Job Order

410 411
Direct materials cost $ 9,100 $ 59,400
Direct labour cost $ 600 $ 11,950
Number of direct labour hours 30 375
Number of parts 500 2,500
Number of machine hours 200 1,100
Number of job orders shipped 1 1
Number of units 10 240
Number of hours in assembly 3 40

Suppose the company had not adopted an ABC system. Compute the manufacturing cost per unit for job orders 410 and 411 under the old, traditional costing system. (Round answers to 2 decimal places, e.g. 255.66.)

Job Order

410 411
Manufacturing cost per unit $

$

     

Under the new ABC system, compute the manufacturing cost per unit for job orders 410 and 411. (Round answers to 0 decimal places, e.g. 255.)

Job Order

410 411
Manufacturing cost per unit $

  

$
Answer #1
1
Job Order
410 411
Direct materials cost 9100 59400
Direct labour cost 600 11950
Indirect manufacturing costs 3000 37500
Total manufacturing costs 12700 108850
Divide by Units 10 240
Manufacturing cost per unit 1270.00 453.54
2
Job Order
410 411
Direct materials cost 9100 59400
Direct labour cost 600 11950
Indirect manufacturing costs:
Material handling 200 1000
Milling 7200 39600
Grinding 350 1750
Assembly 15 200
Inspection 300 7200
Shipping 1600 1600
Total manufacturing costs 19365 122700
Divide by Units 10 240
Manufacturing cost per unit 1936.50 511.25
Workings:
Job Order
410 411
Indirect manufacturing costs:
Material handling =0.4*500 =0.4*2500
Milling =36*200 =36*1100
Grinding =0.7*500 =0.7*2500
Assembly =5*3 =5*40
Inspection =30*10 =30*240
Shipping =1600*1 =1600*1